OpenStudy (anonymous):

A mixture of 40 kg of bird's seed is worth 6 pesos per kg to make up by taking some worth 4.50 pesos per kg and some worth 8.50 per kg. How many kg of each should be taken? Answer please...solution

OpenStudy (dumbcow):

A+B = 40 4.5A + 8.5B = 40*6 = 240 use substitution A = 40-B 4.5(40-B) +8.5B = 240 180 -4.5B +8.5B = 240 4B = 60 B = 15 A = 40-15 = 25

OpenStudy (anonymous):

what is the final answer?

OpenStudy (dumbcow):

what? are you having difficulty following my work? A is amount of seed worth 4.5 pesos B is amount of seed worth 8.5 pesos i solved for A and B

OpenStudy (anonymous):

can you change the variable into x and y??

OpenStudy (dumbcow):

sure, you can make the variables whatever make sense to you

OpenStudy (dumbcow):

so i call it A, you can call it x, someone else can call it P but either way it still represents the same thing which is the number of kg of bird seed worth 4.5 pesos